OpenStudy (pottersheep):

Pleasee help me. How do I plug in 14meters into this equation to find the time?

OpenStudy (pottersheep):

D = depth D(t) = 5sin30(t-5) + 16. The depth is 14...I'm supposed to find time. (t).

OpenStudy (pottersheep):

? ~

OpenStudy (mathmate):

So you're solving 14=5sin(30(t-5)) + 16?

OpenStudy (pottersheep):

yep

OpenStudy (mathmate):

What if we transpose 16 to the other side and continue from: 5sin(30(t-5)) = -2

OpenStudy (pottersheep):

true

OpenStudy (mathmate):

What do you get for solution of sin(30(t-5))=-2/5

OpenStudy (pottersheep):

not sure dont I need t?

OpenStudy (mathmate):

Solve first for the general solution for sin(30(t-5))=-2/5 then t can be solved from the general solution algebraically.

OpenStudy (pottersheep):

thanks for your help......but what do you mean by general solution?

OpenStudy (mathmate):

sin is a periodic function, so sin(x)=1/2 has a solution of pi/6, or 5pi/6, or pi/6+2pi, ... The general solution is therefore pi/6+2k(pi) or 5pi/6+2k(pi). where k is an integer. Since the question does not specify the limits of t, the solution should be derived from the general solution. I hope I made my point clear.

OpenStudy (pottersheep):

Probably, but that's not the way we've learnt it . Thanks anyways

OpenStudy (mathmate):

You're welcome!

OpenStudy (mathmate):

If you accept a solution within a range, then solve for sin(x)=-2/5 (approx. 2.73 for 0<x<pi). since 30(t-5)=2.73 (assuming you are working with radians) you can solve for t (for 0<x<2pi)

OpenStudy (anonymous):

\[t=\frac{1}{30} \left(150-\sin ^{-1}\left(\frac{2}{5}\right)\right) \]A detailed solution is attached.